# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v2

@@ Abstract
Recent advances in learning-based control have increased interest in stable inversion to meet growing performance demands. Here, we establish necessary and sufficient conditions for stable inversion, addressing challenges in non-minimum phase, non-square, and singular systems. An $\mathcal{H}_{\infty}$-based algebraic approximation is introduced for near-perfect tracking without preview. Additionally, we propose a novel robust control strategy combining the nominal model with dual feedforward control to form a feedback structure. Numerical comparison demonstrates the approach’s effectiveness.
@@ I Introduction
Recent advancements include stable inversion for SISO affine systems [8], square systems in continuous time [9] and discrete time [10], and approximate inversion [11]. Additionally, strong inversion for handling initial states in multivariable systems [12] and geometric methods for convolution-based inversion [13] have further expanded the field.
As the field evolves, the need for system classification and corresponding solutions becomes evident. The primary objective of inversion-based approaches is to determine a bounded input that accurately reproduces the desired or given output. However, this goal poses challenges, particularly with non-minimum phase systems, which complicate the establishment of stable inputs under causality without infinite pre-actuation [7]. Furthermore, stable inversion can impact tracking performance and stability in the presence of uncertainties [14]. Merging learning-based control with stable inversion partially addresses these issues [8, 15].
@@
In this paper, we extend the concept of Zero Phase Error Tracking (ZPET) [6] to continuous-time systems. In this setting, we may apply ZPET for compensating the phase shifts introduced by unstable zeros, but ZPET compensation alone does not make sense due to the high-pass nature of these zeros. However, if the overall compensator is designed to suppress these high-pass effects, ZPET remains effective at low frequencies, enabling near-perfect tracking of low-frequency sinusoidal references. The main contributions are:
@@
3) Our approach guarantees robustness without learning mechanisms. When limitations or uncertainties make the output set only partially reachable, we prove that the tracking error converges to an inevitable yet a bounded residual.
@@
We organize this paper as follow: Section II provides the preliminaries. Section III presents the main results. Section IV illustrates numerical examples. Section V concludes the paper.
@@ II-A Algebraic Preliminaries
Let $\mathbb{R}$ be a field of real numbers. Consider the set of all polynomials in $s$ with coefficients in $\mathbb{R}$. This set forms a commutative ring over $\mathbb{R}$ and is denoted by $\mathbb{R}[s]$. If this ring, say $\mathcal{R}$, has an identity element and no zero divisors, then $\mathcal{R}$ is an integral domain. Note that $\mathbb{R}[s]$ is also a Euclidean Domain. Moreover, the set of all such rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [20], and $\mathbb{R}[s]\subset\mathbb{R}(s)$ holds. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$, $\mathbb{R}[s]$, and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$, $\mathbb{R}[s]^{n_{y}\times n_{u}}$, and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively. Any $n_{y}\times n_{u}$ matrix, say $\grave{P}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, over $\mathcal{R}$ can be factorized (see [21, Theorem 2.1]) as
@@
Suppose $\text{rank}_{\mathbb{R}(s)}(P)=r$, where $1\leq r\leq\min(n_{u},n_{y})$. Define $J=\{j_{1},\dots,j_{r}\}\subseteq\{1,\dots,n\}$ as an ordered indexed set corresponding to $P(s)$’s linearly independent columns. We then define the matrix $L$ as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi(u,0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [25, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and casual ($t>0$).
@@
Under Assumption 1, let $\Xi$ denotes the forward system given in (6), and let $y_{1}(t)$ be an observed output trajectory of $\Xi$. We define the right inverse $\Xi_{i}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, which, given $y_{1}(t)$ and a prescribed inverse system’s initial states, $\bar{x}_{0}\in\mathbb{X}$, produces the control input
$u_{inv}(t)=\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty$ (7)
@@
For approximate inversion, among all possible approximations, our goal is to find $\Xi_{i}^{\Gamma}$ that minimizes $e(t)$ and satisfies
@@
The terminology in Problem 1, adopted from [26], allows $y_{1}(t)$ to represent desired trajectories, as in [10]. Assumption 1 and $\|e(t)\|_{\infty}<\infty$ ensure Lyapunov stability for the error dynamics in (8). While our objective of minimizing $e(t)$ is in line with [11], we further introduce a residual term $\varphi$ in (9) capturing the inevitable error due to system limitations.
@@ III-A Right Inverse
Assumption 1 and derivation of (10) show that all rational function entries of $P(s)$ and $Y(s)$ in (11) shares the same greatest common denominator, $\Delta=\det(sI-A)\in\mathbb{R}[s]$. To cancel out $1/\Delta$, we multiply (11) by $\Delta$, yielding
@@
where $\grave{P}(s)\triangleq U_{R}(s)S_{m}(s)V_{R}(s)U(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, $U(s)\in\mathbb{R}(s)^{n_{u}}$, and $\grave{Y}(s)\in\mathbb{R}[s]^{n_{y}}$. Then, using the invariant factor theorem given by (II-A) on $\grave{P}(s)$ in (13), we have
@@ III-B Algebraic Necessary and Sufficient Conditions
Consider $\Xi$ in (6) with P1-P2 and $\Xi^{\Gamma}_{i}$ in (III-A). Then sufficient and necessary condition for Problem 1.a. is
@@
Note that $e(t)=0$ implies (11) and (13) hold and vice-verca. We can now proceed to the proof for both directions.
@@
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}^{\Gamma}_{R1}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}^{\Gamma}_{R1}$ for condition C2. Here, the transformation $L_{T}(s)\in\mathbb{R}(s)^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with $\det(L_{T})\neq 0$ (see [20, Section 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
where the equality and solution are preserved. Now we can re-write $L_{T}(s)$ as $L_{T}(s)=\frac{1}{\Delta_{L}}L^{\prime}_{T}$ s.t. $L^{\prime}_{T}\in\mathbb{R}[s]^{n_{y}\times n_{y}}$, and the scalar $\Delta_{L}\in\mathbb{R}[s]$. So, multiplying (17) by $\Delta_{L}$ to yields
$\grave{P}_{m}(s)\Delta_{L}U(s)=L^{\prime}_{T}(s)\grave{Y}(s)$ (18)
[…]
$\underbrace{y^{*}_{z}\grave{P}_{m}(z)}_{=0}\Delta_{L}(z)U(z)=\underbrace{y^{*}_{z}L^{\prime}_{T}(z)}_{\neq 0}\grave{Y}(z)$ (19)
Since the left-hand side of (19) becomes zero and since $\det(L_{T})\neq 0$ and $1/\Delta_{L}(s)$ is stable, the only way to satisfy the equality is $\grave{Y}(z)=0$.
@@
(22) and (23) are leading to a contradiction. Thus $U(s)$ includes only LHP poles. The proof is now complete. ∎
@@
For the counterpart involving left inverses in contexts such as fault detection, see [12, 26], which provide necessary and sufficient conditions for invertibility under the assumptions of full-rank $P(s)$, $D$ and with $Z_{R}=\emptyset$.
@@ III-C Almost Necessary and Sufficient Conditions
By adhering to analytical constraints (e.g., the waterbed effect, logarithmic integrals, and interpolation bounds) defined in [30] and following the procedure in [31], we ensure $\|\mathcal{F}_{l}(G,K)\|_{\infty}=\gamma<\infty$.
By letting $\epsilon_{i}\to 0$, we replace all approximate integrators in $K$ with pure integrators. Moreover If $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))<\mathbb{E}[N^{2}]$, then $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))=0$ where $\mathbb{E}[N^{2}]$ denotes the expected value mean square of random noise in dB .
@@
Now let us define $P_{p}(s)\in\mathbb{R}(s)$ which is a scalar approximation of MIMO $P(s)$ in (11). To do it first consider one entry of $P(s)$ for $n_{u}<n_{y}$ (taking an average of an entire column is also an option). Because of Assumption 1, the characteristic equation of $P_{p}(s)$, $\Delta_{p}$, equals $\Delta$. Then, if for any RHP zeros of $P(s)$ does not a RHP zeros of $P_{p}(s)$, zero augmentation as $P_{p}(s)=z^{+}(s)P_{p}(s)$ is employed. Then, substitute $P$ with $P_{p}$ in (25) and by letting $\epsilon_{i}\to 0$ solve (III-C) to get stabilizing $K_{p}(s)\in\mathbb{R}(s)$, virtual loop’s control system over the scalar approximation $P_{P}(s)$. Then the scalar complementary sensitivity function for this modified virtual loop is given by
@@
The next corollary provides another approximate inverse, derived from Theorem 1, that also works even with singular systems but allows for less design freedom. To achieve this, one may consider a scalar $z^{+}/z_{d}^{+}$ for the output redefinition, where
@@
Theorems 3, 4 and Corollary 3 each provide a stable approximation when $Y(z)\neq 0$ for any $z\in Z_{R}$. Note also that the control inputs $u_{inv}(t)$ in (28)–(34) are not unique. In Theorem 3, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ always holds because $P(s)$ is square/overactuated and full rank. However, Theorem 4 and Corollary 3 permit $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$, leading to $\varphi\neq 0$ in (9).
All results assume a nominal system $P(s)$, but modeling errors, numerical problems, and unknown forward system initial states can lead instabilities. These implications, along with structures over $\varphi$, will be discussed in the next section.

### v3

@@ Abstract
In this paper, we establish necessary and sufficient conditions for stable inversion, addressing challenges in non-minimum phase, non-square, and singular systems. An $\mathcal{H}_{\infty}$-based algebraic approximation is introduced for near-perfect tracking without preview. Additionally, we propose a novel robust control strategy combining the nominal model with dual feedforward control to form a feedback structure. Numerical comparison demonstrates the approach’s effectiveness.
@@ I Introduction
Recent advancements include stable inversion for SISO affine systems [8], square systems in continuous time [9] and discrete time [10], and approximate inversion with preview[11]. Additionally, strong inversion for handling initial states in multivariable systems [12] and geometric methods for convolution-based inversion [13] have further expanded the field.
As the field evolves, the need for system classification and corresponding solutions becomes evident. The primary objective of inversion-based approaches is to determine a bounded input that accurately reproduces the desired or given output. However, this goal poses challenges, particularly with non-minimum phase systems, which complicate the establishment of stable inputs under causality without infinite pre-actuation. Furthermore, stable inversion can impact tracking performance and stability in the presence of uncertainties [14], where merging learning-based control with stable inversion partially addresses these issues [8, 15].
@@
In this paper, we extend the concept of Zero Phase Error Tracking (ZPET) [6] to continuous-time multivariable systems. In continuous time, we may apply ZPET for compensating the phase shifts introduced by unstable zeros, but ZPET compensation alone does not make sense due to the high-pass nature of these zeros. However, if the overall compensator is designed to suppress these high-pass effects, ZPET remains effective at low frequencies. This insight motivates three contributions:
@@
3) We guarantee robustness without learning-based mechanisms. When limitations or uncertainties make the output set only partially reachable, we prove that the tracking error converges to an inevitable yet bounded residual.
@@
We organize this paper as follows: Section II provides the preliminaries. Section III presents the main results. Section IV illustrates numerical examples. Section V concludes the paper.
@@ II-A Algebraic Preliminaries
Let $\mathbb{R}$ be a field of real numbers. Consider the set of all polynomials in $s$ with coefficients in $\mathbb{R}$. This set forms a commutative ring over $\mathbb{R}$ and is denoted by $\mathbb{R}[s]$. If this ring, say $\mathcal{R}$, has an identity element and no zero divisors, then $\mathcal{R}$ is an integral domain. Note that $\mathbb{R}[s]$ is also a Euclidean Domain. Moreover, the set of all such rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [20], and $\mathbb{R}[s]\subset\mathbb{R}(s)$ holds. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$, $\mathbb{R}[s]$, and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$, $\mathbb{R}[s]^{n_{y}\times n_{u}}$, and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively. Any $n_{y}\times n_{u}$ matrix, say $\grave{P}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, over $\mathcal{R}$ can be factorized (see Invariant Factor Theorem [21, Theorem 2.1]) as
@@
Suppose $\text{rank}_{\mathbb{R}(s)}(P)=r$, where $1\leq r\leq\min(n_{u},n_{y})$. Define $J=\{j_{1},\dots,j_{r}\}\subseteq\{1,\dots,n_{\color[rgb]{0,0,0}u}\}$ as an ordered indexed set corresponding to $P(s)$’s linearly independent columns. We then define the matrix $L$ as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi(u,0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [25, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and causal ($t>0$).
@@
Under Assumption 1, let $\Xi$ denote the forward system given in (6), and let $y_{1}(t)$ be an observed output trajectory. We define the right inverse $\Xi_{i}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, which, given $y_{1}(t)$ and a prescribed inverse system’s initial states, $\bar{x}_{0}\in\mathbb{X}$, produces the control input
$u_{inv}(t)=\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty{\color[rgb]{0,0,0}.}$ (7)
@@
Here, our goal is to construct a stabilizing inverse operator $\Xi_{i}^{\Gamma}$such that the error satisfies
@@
The terminology here (cf. [26]) also lets $y_{1}(t)$ denote a desired trajectory [10]. With Assumption 1 and $\|e\|_{\infty}<\infty$, (8) is Lyapunov-stable. If exact inversion ($\alpha=\varphi=0$) is unattainable, we can employ a robust stabilising approximate inverse—one viable choice among the many approximations acknowledged in [11]—with an irreducible error $\varphi$.
@@ III-A Right Inverse
Under Assumption 1, (10) shows that all rational function entries of $P(s)$ and $Y(s)$ in (11) share the same greatest common denominator, $\Delta=\det(sI-A)\in\mathbb{R}[s]$. To cancel out $1/\Delta$, we multiply (11) by $\Delta$, yielding
@@
where $U(s)\in\mathbb{R}(s)^{n_{u}}$, and $\grave{Y}(s)\in\mathbb{R}{\color[rgb]{0,0,0}(}s{\color[rgb]{0,0,0})}^{n_{y}}$. Then, using the invariant factor theorem given by (II-A) on $\grave{P}(s)$ in (13), we have
@@ III-B Algebraic Necessary and Sufficient Conditions
Consider $\Xi$ in (6) with P1-P2 and $\Xi^{\Gamma}_{i}$ in (III-A). Then necessary and sufficient conditions for Problem 1.a. are
@@
Note that $e(t)=0$ implies (11) and (13) hold and vice versa. We can now proceed to the proof for both directions.
@@
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}_{R1}^{\color[rgb]{0,0,0}\ell}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}_{R1}^{\color[rgb]{0,0,0}\ell}$ for condition C2. Here, the $\mathcal{R}$-Unimodular transformation $L_{T}(s)\in\mathbb{R}{\color[rgb]{0,0,0}[}s{\color[rgb]{0,0,0}]}^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with constant (degree zero) determinant (see [20, Remark 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
Now define ${\color[rgb]{0,0,0}\grave{Y}}(s)\triangleq\frac{1}{\Delta_{\color[rgb]{0,0,0}Y}}{{\color[rgb]{0,0,0}\grave{Y}_{N}}}(s)$ such that ${\color[rgb]{0,0,0}\grave{Y}_{N}}(s)\in\mathbb{R}[s]^{n_{y}}$, and the scalar $\Delta_{\color[rgb]{0,0,0}Y}\in\mathbb{R}[s]$. So, multiplying (17) by $\Delta_{\color[rgb]{0,0,0}Y}$ yields
$\grave{P}_{m}(s)\Delta_{\color[rgb]{0,0,0}Y}U(s)=L_{T}(s)\grave{Y}_{\color[rgb]{0,0,0}N}(s)$ (18)
Note that $\grave{P}_{m}(s)$ is a minimal polynomial basis and the right-hand side of (18) is pure polynomial, so based on Adapted Forney’s Theorem, $\Delta_{\color[rgb]{0,0,0}Y}U(s)$ must be polynomial. It means that $\Delta_{\color[rgb]{0,0,0}Y}$ captures all poles of $U(s)$. Moreover, non-singular transformations of $\grave{P}(s)$ preserves the invariant zeros [28] so that $y^{*}_{z}\grave{P}_{m}(z)=0u^{*}_{z}=0$. Left multiplying $y^{*}_{z}$ to (18) yields
$\underbrace{y^{*}_{z}\grave{P}_{m}(z)}_{=0}\Delta_{\color[rgb]{0,0,0}Y}(z)U(z)=\underbrace{y^{*}_{z}L_{T}(z)}_{\neq 0}\grave{Y}_{\color[rgb]{0,0,0}N}(z)$ (19)
Since the left-hand side of (19) becomes zero, $\det(L_{T}{\color[rgb]{0,0,0}}){\color[rgb]{0,0,0}=\text{constant}}\neq 0$ (by polynomial-unimodularity), and $1/\Delta_{\color[rgb]{0,0,0}Y}$ is stable because $U(s)\in\mathcal{RH}_{\infty}$ is assumed in this direction, the only way to satisfy the equality is $\grave{Y}_{\color[rgb]{0,0,0}N}(z)=0$ implying $Y(z)=0$.
@@
(22) and (23) lead to a contradiction. Thus $U(s)$ includes only LHP poles. The proof is now complete. ∎
@@
For the counterpart involving left inverses in contexts such as fault detection, see [12, 26]. These works give necessary and sufficient conditions for invertibility under the assumptions of full-rank $P(s)$, $D$ and with $Z_{R}=\emptyset$.
@@ III-C Almost Necessary and Sufficient Conditions
Respecting the analytic limits (waterbed/Bode, logarithmic-integral and interpolation bounds) in [30] and leveraging the Youla-based convexification of (III-C) described in [31, Sec. 3.3], we ensure $\|\mathcal{F}_{l}(G,K)\|_{\infty}=\gamma<\infty.$
By letting $\epsilon_{i}\to 0$, we can replace all approximate integrators in $K$ with pure integrators. Moreover If $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))<\mathbb{E}[N^{2}]$, then $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))=0$ where $\mathbb{E}[N^{2}]$ denotes the expected value mean square of random noise in dB .
@@
Now let us define $P_{p}(s)\in\mathbb{R}(s)$ which is a scalar approximation of MIMO $P(s)$ in (11). To do it first consider one entry of $P(s)$ for $n_{u}<n_{y}$ (taking an average of an entire column is also an option). Because of Assumption 1, the characteristic equation of $P_{p}(s)$, $\Delta_{p}$, equals $\Delta$. Then, if for any RHP zeros of $P(s)$ does not have an RHP zero of $P_{p}(s)$, zero augmentation as $P_{p}(s)=z^{+}(s)P_{p}(s)$ is employed. Then, substitute $P$ with $P_{p}$ in (25) and by letting $\epsilon_{i}\to 0$ solve (III-C) to get stabilizing $K_{p}(s)\in\mathbb{R}(s)$, virtual loop’s control system over the scalar approximation $P_{\color[rgb]{0,0,0}p}(s)$. Then the scalar complementary sensitivity function for this modified virtual loop is given by
@@
The next corollary provides another approximate inverse, derived from Theorem 1, that accommodates singular systems. To achieve this, one may, for instance, introduce sufficient low-pass characteristics, or alternatively apply a scalar output redefinition $z^{+}/z_{d}^{+}$, where
@@
Theorems 3, 4, and Corollary 3 each provide a stable approximation when $Y(z)\neq 0$ for all $z\in Z_{R}$. Note also that the control inputs $u_{\text{inv}}(t)$ defined by (28), (34), and (38) are not unique. In Theorem 3, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ always holds because $P(s)$ is either square or overactuated and has full column rank. However, Theorem 4 and Corollary 3 allow for $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$, resulting in $\varphi\neq 0$ in (9).
All results thus far assume a nominal system $P(s)$, but modeling errors, numerical problems, and unknown forward system initial states can lead to instabilities. These implications, along with structures over $\varphi$, will be discussed in the next section.